99 is 180% of what number?

Answer by Maths68(1474)

Let 99 be the 180% of x
x(180%)=99
x(180/100)=99
180x/100=99
180x=99*100
180x=9900
180x/180=9900/180
x=990/18
x=55

99 is 180% of 55
Check
=====
1% of 55 = 55/100
1% of 55 = 0.55
180% of 55 = 0.55*180
180% of 55 = 99